Problem 15-4A (Part Level Submission) Agassi Company uses a job order cost system in each of its three manufacturing departments. Manufacturing overhead is applied to jobs on the basis of direct labor cost in Department D, direct labor hours in Department E, and machine hours in Department K. In establishing the predetermined overhead rates for 2017, the following estimates were made for the year. Department D E K Manufacturing overhead $854,000 $1,500,000 $1,020,000 Direct labor costs $1,220,000 $1,875,000 $675,000 Direct labor hours 150,000 125,000 60,000 Machine hours 600,000 750,000 120,000 During January, the job cost sheets showed the following costs and production data. Department D E K Direct materials used $210,000 $189,000 $117,000 Direct labor costs $180,000 $165,000 $56,250 Manufacturing overhead incurred $148,500 $186,000 $118,500 Direct labor hours 12,000 16,500 5,250 Machine hours 51,000 67,500 10,390 (a) Your answer is correct. Compute the predetermined overhead rate for each department. (Round answers to 2 decimal places, e.g. 10.25.) Overhead rate Department D % Department E $ per direct labor hour Department K $ per machine hour Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Attempts: 1 of 3 used (b) Compute the total manufacturing costs assigned to jobs in January in each department. (Round answers to 0 decimal places, e.g. 2,500.) Manufacturing Costs Department D $ Department E $ Department K $
Answer #1
Predetermined
overhead rate
Department D 70% =854000/1220000
Department E 12.00 =1500000/125000
Department K 8.50 =1020000/120000
Department D Department E Department K
Direct materials used 210000 189000 117000
Direct labor costs 180000 165000 56250
Manufacturing overhead applied 126000 198000 88315
Manufacturing Costs 516000 552000 261565
Workings:
Manufacturing overhead applied:
Department D Department E Department K
=180000*70% =16500*12 =10390*8.5